At the beginning of 20X1, the accounting records of Friends Corp. reported the following:

Preferred shares, 7,200 shares outstanding, no-par $ 238,320
Common shares, 181,000 shares outstanding, no-par 539,380
Contributed capital on common share retirement 110,900
Retained earnings 554,500


During the year, the company acquired and retired shares, while other shares were issued:

15 March 25,100 common shares bought and retired at $4 per share
16 March 4,000 preferred shares bought and retired at $35.50 per share
20 May 9,100 common shares bought and retired at $1 per share
25 May 1,500 preferred shares bought and retired at $18.40 per share
30 May 10,700 common shares issued at $13.40 per share
15 Nov. 4,800 common shares bought and retired at $28 per share


Required:
1. Give journal entries to record each share retirement transaction. (Round intermediate calculations to 2 decimal places and Round your final answers to the nearest whole dollar. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)




2. Calculate the closing balance in each account in shareholders’ equity. (Round intermediate calculations to 2 decimal places. Round your final answers to the nearest whole dollar.)
Answer #1

1.

Journal Entries
Date Accounts Debit ($) Credit ($)
15 March Common Shares Dr 74798 =539380/181000*25100
Contributed capital on common share retirement Dr 25602
Cash 100400 =25100*4
(Being 25,100 common shares bought and retired at $4 per share.)
16 March Preference Shares Dr 132400 =238320/7200*4000
Contributed capital on Preference share retirement Dr 9600
Cash 142000 =4000*35.5
(Being 4,000 preferred shares bought and retired at $35.50 per share.)
20 May Common Shares Dr 27118 =539380/181000*9100
Cash 9100 =9100*1
Contributed capital on common share retirement 18018
(Being 9,100 common shares bought and retired at $1 per share.)
25 May Preference Shares Dr 49650 =238320/7200*1500
Cash 27600 =1500*18.4
Contributed capital on Preference share retirement 22050
(Being 1,500 preferred shares bought and retired at $18.40 per share.)
30 May Cash Dr 143380 =10700*13.4
To Common Shares 143380
(Being 10,700 common shares issued at $13.40 per share.)
15 Nov Common Shares Dr 17701.91 =(539380-74798-27118+143380)/(181000-25100-9100+10700)*4800
Contributed capital on common share retirement Dr 103316 Balance
Retained Earnings Dr 13382.1 Remaining from retained earnings
To Cash 134400 =4800*28
(Being 4,800 common shares bought and retired at $28 per share.)

2.

Shareholders' Equity
Preferred shares, 1700 shares outstanding, no-par $56,270.0 =238320-132400-49650
Common shares, 152,700 shares outstanding, no-par $5,63,142.1 =539380-74798-27118+143380-17701.91
Contributed capital on common share retirement $0.0 =110900-25602+18018-103316
Contributed capital on Preference share retirement $12,450.0 =-9600+22050
Retained earnings $5,41,117.9 =554500-13382.1
Total Shareholders' Equity $11,72,980.0